2 Problem 14-15 Early extinguishment; effective interest [L014-5] 2.5 of Twin Digital Corporations balance sheet as of December 31, 2017, included 10% bonds million and a remaining discount of $1 milion. Disclosure notes indicate the bonds were issued to Interest expense Is recorded at the effective interest rate and paid on January 1 and July 1 of each year. On July 1, 2018, Twin Digital retired the bonds at 103 ($51.5 million) before their scheduled maturity. Print Required: References 1.&2.Prepare the necessary journal entries for Twin Digital on July 1, 2018. (If no entry is required for a transaction/event, select No journal entry required in the first account field. Enter your answers in whole dollar.) No Date July 01, 2018 Interest expense 440.000 K Prex 2 of 2E Next
Answer #1
Debit Credit
July 01, 2018 Interest expense 2940000 =(50000000-1000000)*12%/2
        Cash 2500000 =50000000*10%/2
        Discount on bonds payable 440000
July 01, 2018 Bonds payable 50000000
Loss on early extinguishment 2060000
         Cash 51500000
         Discount on bonds payable 560000 =1000000-440000